OpenStudy (anonymous):

What is the slope of a line that passes through the point (−5, 3) and is parallel to a line that passes through (2, 13) and (−4, −11)?

OpenStudy (saifoo.khan):

Aroubaaa will Rescue the Moon.

OpenStudy (anonymous):

Find the slope between points (2, 13) and (-4, -11). Because parallel lines have the same slope, that slope will also be the slope of the line that passes through (-5, 3). Use y=m(x-x1)+y1. m= slope, (x1, y1) = (-5, 3)

OpenStudy (anonymous):

4

OpenStudy (aroub):

-11-13/ -4-2 -24/-6 4 y=mx+b , m=4 , ( -5,3) 3=4(-5)+b 3=-9+b 12=b y=4x+12

OpenStudy (anonymous):

the equation is y=4x + 23

OpenStudy (aroub):

oh yes sry i guess i made something wrong..!! :(